# A Reverse Mamba Attention Network for Pathological Liver Segmentation

### v1

@@ Abstract
We present RMA-Mamba, a novel architecture that advances the capabilities of vision state space models through a specialized reverse mamba attention module (RMA). The key innovation lies in RMA-Mamba’s ability to capture long-range dependencies while maintaining precise local feature representation through its hierarchical processing pipeline. By integrating Vision Mamba (VMamba)’s efficient sequence modeling with RMA’s targeted feature refinement, our architecture achieves superior feature learning across multiple scales. This dual-mechanism approach enables robust handling of complex morphological patterns while maintaining computational efficiency. We demonstrate RMA-Mamba’s effectiveness in the challenging domain of pathological liver segmentation (from both CT and MRI), where traditional segmentation approaches often fail due to tissue variations. When evaluated on a newly introduced cirrhotic liver dataset (CirrMRI600+) of T2-weighted MRI scans, RMA-Mamba achieves the state-of-the-art performance with a Dice coefficient of 92.08%, mean IoU of 87.36%, and recall of 92.96%. The architecture’s generalizability is further validated on the cancerous liver segmentation from CT scans (LiTS: Liver Tumor Segmentation dataset), yielding a Dice score of 92.9% and mIoU of 88.99%. The source code of the proposed RMA-Mamba is available at https://github.com/JunZengz/RMAMamba.
Keywords:
Mamba Reverse Attention Liver segmentation Cirrhosis Liver Cancer.
1 Introduction
Technical motivation. The Transformer architecture [37] significantly advanced sequence modeling by employing self-attention mechanisms to capture long-range dependencies without relying on recurrence. Despite their success, Transformers face limitations in handling long sequences due to the quadratic complexity of the self-attention operation with respect to sequence length. To address this challenge, recent studies have explored the integration of State Space Models (SSMs) into Transformer architectures, leading to the development of Mamba-style architectures. These architectures converge the strengths of Transformers and SSMs by replacing the standard self-attention with linear recurrent layers derived from state space formulations. Specifically, they utilize parameterizations that enable efficient computation of long-range dependencies with linear complexity, such as the Structured State Space Sequence model (S4). This convergence allows Mamba architectures to maintain the expressive power of Transformers while significantly improving scalability and efficiency in modeling long sequences. The evolution from pure self-attention mechanisms to incorporating SSMs represents a significant step forward in designing architectures that are both computationally efficient and capable of capturing complex dependencies in high-resolution computer vision tasks [9, 29, 39].
@@ 1 Introduction
Clinical motivation. The liver, vital for metabolism and digestion, is increasingly affected by cirrhosis [31] and represents a critical site for both primary and metastatic cancer development. As the sixth most common location for primary malignancies, it also frequently harbors metastases from other abdominal organs, particularly the colon and pancreas [4, 1]. Early detection of liver pathologies significantly impacts patient survival rates [41] and treatment options. While modern imaging modalities like computed tomography (CT) and magnetic resonance imaging (MRI) provide detailed visualization of hepatic tissue, the current clinical workflow (still!) relies heavily on manual or semi-automated slice-by-slice contour delineation by radiologists. This process is not only time-intensive but also subject to inter-observer variability and expertise-dependent accuracy. The development of robust computer-assisted segmentation algorithms for pathological liver tissue addresses these clinical challenges by offering consistent, automated analysis of medical images. Such tools promise to reduce manual annotation burden while potentially improving diagnostic accuracy through standardized quantitative assessment.
While liver segmentation has seen remarkable progress over recent decades  [3, 22, 46, 18, 19, 8], the specific challenge of cirrhotic liver segmentation from MRI remains underexplored, and cancerous liver segmentation (from CT and MRI) have significant limitations. This gap is particularly concerning given that early-stage cirrhosis often goes undetected in clinical practice, leading to disease progression toward decompensated cirrhosis or hepatocellular carcinoma (HCC). The increasing mortality rates associated with this progression underscore the urgent need for robust segmentation tools that can later be used in a computer aided diagnosis and quantification classifier to reliably identify cirrhotic changes across diverse patient populations and imaging protocols. The recent introduction of the CirrMRI600+ dataset [14] represents a significant step toward addressing this unmet clinical need. This comprehensive dataset enables the development and validation of advanced segmentation algorithms capable of handling the subtle tissue changes characteristic of early cirrhosis, while maintaining robustness across different MRI scanners and patient demographics.
@@
  • Rich benchmarking results: Extensive benchmarking against thirteen state-of-the-art methods demonstrates RMA-Mamba’s superior performance, achieving a dice coefficient of 92.08%, mean IoU of 87.36%, and recall of 92.96% on CirrMRI600+. These results validate our architectural innovations and establish new performance standards for cirrhotic and cancerous liver segmentation. The model’s effectiveness is further confirmed through rigorous evaluation on the LiTS dataset, demonstrating robust generalization across different imaging protocols.
2 Related Work
2.1 State Space Model
SSMs have emerged as powerful architectures for sequence modeling [29, 6, 12], with each iteration addressing key computational and performance challenges. The Linear State Space Layer (LSSL) [11] pioneered the use of HIPPO (High-order Polynomial Projection Operators) matrices and structured matrices for capturing long-range dependencies. However, its computational complexity limited practical applications for extended sequences.
@@ 2.1 State Space Model
2.2 Liver segmentation
Deep learning has transformed medical image segmentation, initially with UNet [32], later with nnUnet [47], establishing a foundational encoder-decoder architecture that efficiently combines high-resolution features with contextual information through skip connections. These architectures have spawned numerous innovations in liver segmentation [24, 21, 23, 38, 7], each addressing specific aspects of the segmentation challenge.
@@ 2.2 Liver segmentation
Recent architectures have focused on multi-scale feature learning and boundary refinement. Xie et al. [43] proposed a multi-scale context extraction module with external attention and boundary correction mechanisms. Liu et al. [26] introduced a hybrid attention approach combining global dependencies with local feature focus. Moving beyond traditional convolutional architectures, Jha et al. [19] leveraged transformer-based vision models with residual connections, demonstrating superior computational efficiency and segmentation accuracy.
@@ 3.3 VSS block
3.4 Reverse Mamba Attention Module
The reverse attention mechanism has demonstrated remarkable success in salient object detection [2, 45], particularly in boundary delineation and feature refinement. This success was notably exemplified in PraNet [5], where reverse attention effectively enhanced boundary detection and tissue differentiation, leading to improved segmentation accuracy through targeted error correction. Building on this foundation, RTA-Former [25] advanced the concept by introducing transformer-based reverse attention, demonstrating superior edge segmentation capabilities through sophisticated feature extraction. These successes in boundary refinement and feature enhancement inspired our novel approach: leveraging the computational efficiency of VSS blocks within a reverse attention framework for pathological liver.
@@ 3.5 Loss function
Table 1: Model performance on the CirrMRI600+ dataset.
@@
ResUNet++ [16] IEEE ISM 2019 87.76 82.09 89.75 90.30 88.57 3.68
DoubleU-Net [15] IEEE CBMS 2020 91.30 86.57 91.91 93.30 91.45 3.42
ColonSegNet [13] IEEE Access 2021 87.60 81.67 89.49 90.39 88.30 3.66
NanoNet [17] IEEE CBMS 2021 81.66 74.55 83.50 85.20 82.33 4.12
@@
Table 2: Model performance on the LiTS dataset.
@@
ResUNet++ [16] IEEE ISM 2019 90.94 86.49 90.50 94.72 90.47 2.96
DoubleU-Net [15] IEEE CBMS 2020 92.37 88.34 92.66 94.54 92.29 2.85
ColonSegNet [13] IEEE Access 2021 89.71 84.95 89.81 93.49 89.52 3.03
NanoNet [17] IEEE CBMS 2021 88.33 83.30 88.82 91.84 88.38 3.14
@@
4 Experiments and Results
4.1 Datasets
@@
LiTS [1] is a multicenter dataset collected from various clinical research institutes, containing expert-annotated ground truths for liver and tumor segmentation in 201 abdominal CT scans. The training set includes 130 CT scans with publicly available ground truths, which are preprocessed to generate healthy liver masks. In our experiments, these liver masks were divided into three parts: 11684 slices for the training set, 2745 for the validation set, and 4734 for the test set. The difference in slice counts between the validation and test sets is due to variations in the number of CT slices per patient.
4.2 Implementation
@@
Table 3: Ablation study on the CirrMRI600+ dataset.
@@ 4.3 Evaluation metrics
4.4 Benchmarking
We compared our architectures, RMAMamba-S and RMAMamba-T, with the state-of-the-art deep segmentation methods (U-Net [32], ResUNet++ [16], DoubleU-Net [15], ColonSegNet [13], NanoNet [17], UNext [36]), Transformer-based methods (TransNetR [20], TransResUNet [35], PVTFormer [19]), and SSM-based segmentation methods (VM-UNet [33], VM-UNetV2 [44], UltraLight VM-UNet [42], Mamba-UNet [40]) on the CirrMRI600+ and LiTS datasets. To ensure a fair comparison, all models were trained and evaluated with identical dataset partitioning and hyperparameter settings.
@@ 4.4 Benchmarking
Our extensive experimentation on the pathological liver segmentation benchmark dataset demonstrates RMA-Mamba’s superior performance (Table 1). RMA-Mamba-T outperforms existing CNN and Transformer-based architectures while showing marked improvement over recent SSM-based approaches like VM-UNet and Mamba-UNet. The larger variant, RMAMamba-S, achieves state-of-the-art performance with a Dice score of 92.08%, mIoU of 87.36%, and Hausdorff Distance of 3.39 mm.
@@
Table 2 presents the quantitative results of various state-of-the-art models on the LiTS dataset. We can observe that RMA-Mamba-T achieves exceptional segmentation performance, with a Dice coefficient of 92.94%, mIoU of 88.99%, recall of 92.34%, precision of 95.59%, F2 score of 92.44%, and a HD score of 2.83 mm. RMA-Mamba-T achieved the highest Dice, surpassing PVTFormer by 0.46%, VM-UNetV2 by 0.89%, and ResUNet++ by 2%. Compared to RMA-Mamba-S, RMA-Mamba-T has fewer parameters yet delivers superior performance on the LiTS dataset. Additionally, RMA-Mamba-T offers faster inference speed and lower computational cost. Figure 3 shows that our model provides more accurate segmentation results than other benchmarked state of the art methods.
4.5 Ablation study
Table 3 summarizes the ablation study of the RMA-Mamba on the CirrMRI600+ dataset. As mentioned in Section 3, RMA-Mamba consists of a VMamba backbone, additional VSS blocks beyond the backbone, and RMA modules, where N represents the number of additional VSS blocks. In Table 3, RA denotes the traditional convolution-based reverse attention mechanism, originally introduced in PraNet [5]. The first group (#1) and the second group (#2) represent experimental setups using Vmamba-Tiny and Vmamba-Small as the network backbones, respectively, without the RMA module and additional VSS blocks. The fifth group (#5) and the sixth group (#6) correspond to the RMA-Mamba-T and RMA-Mamba-S architectures, respectively.
@@ 4.5 Ablation study
Table 3 illustrates that both RMA-Mamba-S (#6) and RMA-Mamba-T (#5) demonstrate excellent performance on the CirrMRI600+ dataset. Therefore, we selected #5 as RMA-Mamba-T and #6 as RMA-Mamba-S based on their complementary strengths: RMA-Mamba-T delivers the highest performance despite a larger parameter count, while RMA-Mamba-S maintains strong segmentation accuracy with fewer parameters. Depending on the specific requirements of different tasks, we can flexibly choose between RMA-Mamba-S and RMA-Mamba-T.
5 Conclusion

### v2

@@ Abstract
We present RMA-Mamba, a novel architecture that advances the capabilities of vision state space models through a specialized reverse mamba attention module (RMA). The key innovation lies in RMA-Mamba’s ability to capture long-range dependencies while maintaining precise local feature representation through its hierarchical processing pipeline. By integrating Vision Mamba (VMamba)’s efficient sequence modeling with RMA’s targeted feature refinement, our architecture achieves superior feature learning across multiple scales. This dual-mechanism approach enables robust handling of complex morphological patterns while maintaining computational efficiency. We demonstrate RMA-Mamba’s effectiveness in the challenging domain of pathological liver segmentation (from both CT and MRI), where traditional segmentation approaches often fail due to tissue variations. When evaluated on a newly introduced cirrhotic liver dataset (CirrMRI600+) of T2-weighted MRI scans, RMA-Mamba achieves the state-of-the-art performance with a Dice coefficient of 92.08%, mean IoU of 87.36%, and recall of 92.96%. The architecture’s generalizability is further validated on the cancerous liver segmentation from CT scans (LiTS: Liver Tumor Segmentation dataset), yielding a Dice score of 92.9% and mIoU of 88.99%. Our code is available for public: https://github.com/JunZengz/RMAMamba.
Index Terms:
Vision Mamba, Reverse Attention, Liver segmentation, Cirrhosis, Liver Cancer.
I Introduction
Technical motivation. The Transformer architecture [26] significantly advanced sequence modeling by employing self-attention mechanisms to capture long-range dependencies without relying on recurrence. Despite their success, Transformers face limitations in handling long sequences due to the quadratic complexity of the self-attention operation with respect to sequence length. To address this challenge, recent studies have explored the integration of State Space Models (SSMs) into Transformer architectures, leading to the development of Mamba-style architectures. These architectures converge the strengths of Transformers and SSMs by replacing the standard self-attention with linear recurrent layers derived from state space formulations. Specifically, they utilize parameterizations that enable efficient computation of long-range dependencies with linear complexity, such as the Structured State Space Sequence model (S4). This convergence allows Mamba architectures to maintain the expressive power of Transformers while significantly improving scalability and efficiency in modeling long sequences. The evolution from pure self-attention mechanisms to incorporating SSMs represents a significant step forward in designing architectures that are both computationally efficient and capable of capturing complex dependencies in high-resolution computer vision tasks [7].
@@ I Introduction
Clinical motivation. The liver, vital for metabolism and digestion, is increasingly affected by cirrhosis [20] and represents a critical site for both primary and metastatic cancer development. As the sixth most common location for primary malignancies, it also frequently harbors metastases from other abdominal organs, particularly the colon and pancreas [2]. Early detection of liver pathologies significantly impacts patient survival rates [29] and treatment options. While modern imaging modalities like computed tomography (CT) and magnetic resonance imaging (MRI) provide detailed visualization of hepatic tissue, the current clinical workflow still relies heavily on manual or semi-automated slice-by-slice contour delineation by radiologists. This process is not only time-intensive but also subject to inter-observer variability and expertise-dependent accuracy. The development of robust computer-assisted segmentation algorithms for pathological liver tissue addresses these clinical challenges by offering consistent, automated analysis of medical images. Such tools promise to reduce manual annotation burden while potentially improving diagnostic accuracy through standardized quantitative assessment.
While liver segmentation has seen remarkable progress over recent decades [12, 34, 6], the specific challenge of cirrhotic liver segmentation from MRI remains underexplored, and cancerous liver segmentation (from CT and MRI) have significant limitations. This gap is particularly concerning given that early-stage cirrhosis often goes undetected in clinical practice, leading to disease progression toward decompensated cirrhosis or hepatocellular carcinoma (HCC). The increasing mortality rates associated with this progression underscore the urgent need for robust segmentation tools that can later be used in a computer aided diagnosis and quantification classifier to reliably identify cirrhotic changes across diverse patient populations and imaging protocols. The recent introduction of the CirrMRI600+ dataset [9] represents a significant step toward addressing this unmet clinical need. This comprehensive dataset enables the development and validation of advanced segmentation algorithms capable of handling the subtle tissue changes characteristic of early cirrhosis, while maintaining robustness across different MRI scanners and patient demographics.
@@
  • Rich benchmarking results: Extensive benchmarking against eight state-of-the-art methods demonstrates RMA-Mamba’s superior performance, achieving a dice coefficient of 92.08%, mean IoU of 87.36%, and recall of 92.96% on CirrMRI600+. These results validate our architectural innovations and establish new performance standards for cirrhotic and cancerous liver segmentation. The model’s effectiveness is further confirmed through rigorous evaluation on the LiTS dataset, demonstrating robust generalization across different imaging protocols.
II Related Work
II-A State Space Model
SSMs have emerged as powerful architectures for sequence modeling [4], with each iteration addressing key computational and performance challenges. The Linear State Space Layer (LSSL) [8] pioneered the use of HIPPO (High-order Polynomial Projection Operators) matrices and structured matrices for capturing long-range dependencies. However, its computational complexity limited practical applications for extended sequences.
@@ II-A State Space Model
II-B Liver segmentation
Deep learning has transformed medical image segmentation, initially with UNet [21], later with nnUnet, establishing a foundational encoder-decoder architecture that efficiently combines high-resolution features with contextual information through skip connections. These architectures have spawned numerous innovations in liver segmentation [14, 11, 13, 27, 5], each addressing specific aspects of the segmentation challenge.
@@ II-B Liver segmentation
Recent architectures have focused on multi-scale feature learning and boundary refinement. Xie et al. [31] proposed a multi-scale context extraction module with external attention and boundary correction mechanisms. Liu et al. [16] introduced a hybrid attention approach combining global dependencies with local feature focus.
@@ III-C VSS block
III-D Reverse Mamba Attention Module
The reverse attention mechanism has demonstrated remarkable success in salient object detection [33], particularly in boundary delineation and feature refinement. This success was notably exemplified in PraNet [3], where reverse attention effectively enhanced boundary detection and tissue differentiation, leading to improved segmentation accuracy through targeted error correction. Building on this foundation, RTA-Former [15] advanced the concept by introducing transformer-based reverse attention, demonstrating superior edge segmentation capabilities through sophisticated feature extraction. These successes in boundary refinement and feature enhancement inspired our novel approach: leveraging the computational efficiency of VSS blocks within a reverse attention framework for pathological liver.
@@ III-E Loss function
Table I: Model performance on the CirrMRI600+ dataset.
@@
Table II: Model performance on the LiTS dataset.
@@
IV Experiments and Results
IV-A Datasets
@@
LiTS [1] is a multicenter abdominal CT dataset (201 scans) with expert annotations for liver and tumor segmentation. The training set (130 scans) was used to generate healthy liver masks, which were then divided into training (11684 slices), validation (2745 slices), and test (4734 slices) sets. Slice count variations in validation and test sets reflect patient-specific CT slice numbers.
IV-B Implementation
@@
Table III: Ablation study on the CirrMRI600+ dataset.
@@ IV-C Evaluation metrics
IV-D Benchmarking
We compared our architectures, RMAMamba-S and RMAMamba-T, with the state-of-the-art deep segmentation methods (U-Net [21], UNext [25]), Transformer-based method (TransResUNet [24], TransNetR [10]), and SSM-based segmentation methods (VM-UNet [22], VM-UNetV2 [32], UltraLight VM-UNet [30], Mamba-UNet [28]) on the CirrMRI600+ and LiTS datasets. To ensure a fair comparison, all models were trained and evaluated with identical dataset partitioning and hyperparameter settings.
@@ IV-D Benchmarking
Our extensive experimentation on the pathological liver segmentation benchmark dataset demonstrates RMA-Mamba’s superior performance (Table I). RMA-Mamba-T outperforms existing CNN and Transformer-based architectures while showing marked improvement over recent SSM-based approaches like VM-UNet and Mamba-UNet. The larger variant, RMAMamba-S, achieves state-of-the-art performance with a Dice score of 92.08%, mIoU of 87.36%, and Hausdorff Distance of 3.39 mm.
@@
Table II presents the quantitative results of various state-of-the-art models on the LiTS dataset. We can observe that RMA-Mamba-T achieves exceptional segmentation performance, with a Dice coefficient of 92.94%, mIoU of 88.99%, recall of 92.34%, precision of 95.59%, F2 score of 92.44%, and a HD score of 2.83 mm. RMA-Mamba-T achieved the highest Dice, surpassing PVTFormer by 0.46%, VM-UNetV2 by 0.89%, and ResUNet++ by 2%. Compared to RMA-Mamba-S, RMA-Mamba-T has fewer parameters yet delivers superior performance on the LiTS dataset. Additionally, RMA-Mamba-T offers faster inference speed and lower computational cost. Figure 3 shows that our model provides more accurate segmentation results than other benchmarked state-of-the-art methods.
IV-E Ablation study
Table III summarizes the ablation study of the RMA-Mamba on the CirrMRI600+ dataset. As mentioned in Section III, RMA-Mamba consists of a VMamba backbone, additional VSS blocks beyond the backbone, and RMA modules, where N represents the number of additional VSS blocks. In Table III, RA denotes the traditional convolution-based reverse attention mechanism, originally introduced in PraNet [3]. The first group (#1) and the second group (#2) represent experimental setups using Vmamba-Tiny and Vmamba-Small as the network backbones, respectively, without the RMA module and additional VSS blocks. The fifth group (#5) and the sixth group (#6) correspond to the RMA-Mamba-T and RMA-Mamba-S architectures, respectively.
@@ IV-E Ablation study
Table 3 illustrates that both RMA-Mamba-S (#6) and RMA-Mamba-T (#5) demonstrate excellent performance on the CirrMRI600+ dataset. Therefore, we selected #5 as RMA-Mamba-T and #6 as RMA-Mamba-S for their complementary strengths: RMA-Mamba-T delivers the highest performance despite a larger parameter count, while RMA-Mamba-S maintains strong segmentation accuracy with fewer parameters. Depending on the requirements of different tasks, we can flexibly choose between RMA-Mamba-S and RMA-Mamba-T.
V Conclusion